Suppose, X has the following PDF:
f(x)={Cxe^x/2 // 0 x>0 x<=0
a. Find the value of C.
b. Find the mean of ​X​.
c. Find the 2nd Quartile of ​X​.
d. Also, find the standard deviation of ​X
1
Expert's answer
2020-04-04T16:23:42-0400

"f(x)=\begin{cases}
Cxe^{-x/2}, x>0
\\
0, x\leq 0
\end{cases}"

(here we have "e^{-x\/2}" because "Cx e^{x\/2}" is unbounded, but "\int \limits_0^{\infin}f(x)=1" to be PDF)


a.

"\int \limits _0^{\infin}f(x)dx= \int \limits _0^{\infin}Cxe^{-x/2}dx=\\=-C(2x+4)e^{-x/2} \big|_0^{\infin}=4C \ \Rightarrow C=1/4"

Answer: "C=1\/4"


b.

"\mu =\int \limits_0^{\infin}xf(x)dx =\int \limits_0^{\infin}1/4 x^2e^{-x/2}dx=\\=-1/2(x^2+4x+8)e^{-x/2}\big|_0^{\infin}=4"

Answer: "\mu=4"


c.

"0.5=F(x)=\int \limits _0^xf(t)dt=\\=\int\limits _0^x 1/4te^{-t/2}dt=-1/2(x+2)e^{-x/2}+1,"

"0.5=-1/2(x+2)e^{-x/2}+1, \ \ (x+2)=e^{x/2}, \ \ \ x\approx 3.356"

Answer: the second quartile is "3.356".


d.

"\ \sigma ^2=\int\limits _0^{\infin} (x-\mu)^2f(x)dx= \int\limits _0^{\infin} 1/4(x-4)^2x e^{-x/2}dx=\\=
-1/2(x^3-2x^2+8x+16)e^{-x/2}\big|_0^{\infin}=8,"

"\sigma=\sqrt{\sigma^2}=\sqrt{8}=2\sqrt{2}."

Answer: "\sigma =2\sqrt{2}"